The level of bank deposits is $9,681 and the banks hold $1,046 in required reserves and 765 in excess reserves. If the Fed introduces $100 worth of new reserves worth of new reserves in the economy by how much will the money supply increase.
Answer #1

Required reserves ratio= 1046/9681= 0.108

Excess reserves ratio = 765/9681= 0.079

Money multiplier = 1/(0.108+0.079)= 5.3476

Increase in money supply = 100/5.347= 534.76